OpenStudy (anonymous):

\[Evaluate \int\limits x ^{2}\csc(cosx ^{3})sinx ^{3}dx\]

OpenStudy (anonymous):

use this \[\large \frac{d}{dx} \cos x^3=-3x^2 \sin x^3\]

OpenStudy (anonymous):

so u can start by letting \[\large u= \cos x^3\]

OpenStudy (anonymous):

:) because of integral I forgot how to derive. hehe. lol. :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

anyway thanks.:)